OpenStudy (anonymous):

prove that: \[\frac{ \tan \theta - \cot \theta }{ \tan \theta + \cot \theta } = 2\sin ^{2} \theta\]

OpenStudy (anonymous):

Correcrtion: \[\frac{ \tan \theta - \cot \theta }{ \tan \theta + \cot \theta } = 2\sin ^{2} \theta-1\]

OpenStudy (ajprincess):

Try writing \(\tan\theta\) as \[\frac{\sin\theta}{\cos\theta}\] and \(\cot\theta\) as \[\frac{\cos\theta}{\sin\theta}\] and then simplify.

OpenStudy (ajprincess):

Can u do @Beatles?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

\[\frac{ \sin \theta - \cos \theta }{ \sin \theta + \cos \theta }\]

OpenStudy (anonymous):

this.

zepdrix (zepdrix):

?

zepdrix (zepdrix):

Beatle, when they say "prove", you want to try and alter one side of the equation and make it look like the other. Don't change both sides. :D

OpenStudy (ajprincess):